### v2

@@ 4.4 Variation Graph Construction with LCPan
As illustrated in Figure 5, LCPan applies the following steps for pangenome graph construction:
  1. 1
    Process each input genome up to a specified LCP level to partition into cores.
  2. 2
    Cluster variations that originate within each LCP core.
  3. 3
    Further partition each LCP core to integrate the variations.
  4. 4
    Link the variations to the backbone reference genome.
Finally, for both ends of the sequences, we allocate separate segments to represent the reference sequence completely. Note that there might be at most two such segments in a given ungapped sequence.
@@ Acknowledgments
Adjacency Property: within a substring of length 33, there could be at most 22 characters that could be the starting positions of cores.

### v3

@@ 4.4 Variation Graph Construction with LCPan
Figure 5 illustrates LCPan’s steps for graph construction. Using the selected LCP level, we partition the initial linear genome into LCP cores per chromosome (step 1). The initial linear genome yields a simple linear “backbone” graph where nodes represent cores, and edges are placed between pairs of consecutive cores of the associated LCP level. For each variation in the input VCF file, we apply the respective operation for the variation (step 2) to each of the associated sets of LCP cores in this graph. Each branch created between the consecutive LCP cores generates an alternative path in the graph. If the offset of a variation overflows the current LCP core’s range, we reassign it to the next appropriate LCP core. LCPan then further partitions the core (step 3) to integrate additional variations from alternative haplotypes into the graph (step 4), which, in its final form, is referred to as the variation graph. Finally, for both ends of the sequences, we allocate separate segments to represent the reference sequence completely. Note that there might be at most two such segments in a given ungapped sequence.
The fundamental strength of LCPan lies in the guarantees provided by Locally Consistent Parsing (LCP), particularly regarding partition consistency. This consistency is essential, as it ensures that the partitioned substrings maintain near-uniform lengths. By sequentially processing variation files stored in sorted order, variations within the offsets of LCP cores can be organized and processed concurrently and independently in batches. Moreover, consistent partitioning and labeling facilitate rapid matching, as LCP ensures that sequences are parsed identically and independently, even when separated by megabases. Furthermore, higher LCP levels lead to longer cores and fewer partitions during parsing. Each partition then contains more variations that must be processed together, resulting in more concentrated work per execution unit. Consequently, performance initially improves with the addition of more threads, but after a certain threshold (e.g., 8 threads), the benefit of adding more threads diminishes due to the limits of parallelization.
@@ Acknowledgments
Adjacency Property: within a substring of length $3$, there could be at most $2$ characters that could be the starting positions of cores.